OpenStudy (anonymous):

Identify whether the given value of the variable makes the equation true. 14x = –36; x = –3 A. false B. true

OpenStudy (fibonaccichick666):

do you have any ideas on how to begin?

OpenStudy (anonymous):

i have trouble with the negaitive and positves one

OpenStudy (fibonaccichick666):

? that doesn't really come into play that much here. All you do is plug x in and see if you get the correct answer

OpenStudy (anonymous):

i think your start with 14 - 3-

OpenStudy (anonymous):

plug?

OpenStudy (math&ing001):

@TeenTitansfan1010 it's as easy as replacing x with the given value, like this : 14 * (-3) = -36 and see if it's correct.

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

plug = replace, sorry it's sort of a common term for me

OpenStudy (anonymous):

11? so its false?

OpenStudy (anonymous):

i did 14 X 3-

OpenStudy (anonymous):

= 11

OpenStudy (math&ing001):

@TeenTitansfan1010 you do not subtract, it's a multiplication : 14 * (-3) = - (14*3) = - 42

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

FALSE THEN?

OpenStudy (anonymous):

opps caps

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

TJANKSS!!!!!!!!!

OpenStudy (math&ing001):

Welcome ^^

OpenStudy (anonymous):

THANKS*